# A systematic review of metaheuristics-based and machine learning-driven intrusion detection systems in IoT

### v3

@@ Abstract
Keywords: Internet of Things (IoT), Intrusion Detection Systems (IDS), Machine Learning (ML), Deep Learning (DL), Metaheuristic Algorithms, Cybersecurity, Optimization Techniques.
1 Introduction
@@
The popularity of the Internet of Things devices has spread surprisingly in the last few years. Nowadays, it offers a scalable platform not only for industry, healthcare, and home applications but also for agriculture, vehicular settings, and ultra-sophisticated systems like drone technologies [intro_2_abualigah2021applications]. Alarmingly, this widespread adoption leads to unavoidable security issues as a negative side effect of close connectivity. Transferring susceptible information, such as personal data, patient data, and private business analysis makes such effects more severe and unimaginable. For example, millions of IoT and embedded devices (e.g., DVDs, printers, and IP cameras) were infected by a botnet in 2016, widely known as Mirai botnet[antonakakis2017understanding]. Notably, the attack was spread to several nations and manufacturing organizations and affected around 65 thousand IoT devices within the first 20 hours. Another concerning incident occurred in 2020, where an ADT employee pled guilty to accessing the security cameras of 220 women over 9,600 times during four years [erin]. The defendant routinely added his email address to customers’ “ADT Pulse” accounts and got real-time access to the video feeds from their homes for sexual gratification. In such cases, an IoT-IDS can be used to detect and report on unique visitors, unauthorized access, and malicious activities.
@@
Intrusion detection is one of the most crucial aspects of IoT security. An IDS typically identifies diverse attacks based on predefined rules or specific deviations from normal behavioral patterns. It can identify external and internal attacks on networks or computer systems, surpassing the performance of the traditional firewall. Typically, a firewall works on a set of specific rules, often based on IP addresses, port numbers, and protocols to check which packets are allowed to enter the network. Since firewalls filter packets relying on simple policies, they cannot detect internal or external attacks which require analyzing complex behavioral patterns. On the contrary, an IDS can identify malicious activities by observing the deviations from normal behaviors, which include both simple and complicated patterns. The entire IDS can be divided into two modules: feature engineering (mostly feature selection) and classification or detection. Feature selection aims to select a set of optimal features, discarding the least significant ones to efficiently and faster the classification process and reduce the computational overhead of the system. Feature selection methods can be categorized into three methods: filter, wrapper, and hybrid methods. In the filter method, all features are statistically examined and rated with the help of data (input and target variables). Then based on the rating, less ranked features or features rated below a specific threshold are eliminated before classification. Information gain and linear correlation coefficient [eid2013linear] are well-known metrics used in filter methods. The wrapper techniques outperform filter methods by training and testing a machine-learning model using each subset of features, iteratively. Specifically, these methods rank subsets of features based on their prediction accuracy generated from the machine learning algorithms. However, they are more expensive and time-consuming than the filter approaches. Sequential forward selection, sequential backward selection, stepwise selection [hastie2017extended], hill climbing, etc. are popular wrapper methods. Though filter methods are proven simple, fast, and scalable, they consider feature dependencies to a certain extent, resulting in an inappropriate feature set. Besides, wrapper methods offer better feature sets than filters. But, they become much slower and computationally expensive for a large number of features [wah2018feature].
@@
A group of wrapper methods fall into the category of optimization-based techniques, known as metaheuristics algorithms that overcome the drawbacks of the prior methods. A metaheuristic is a general exploration method that applies to optimize an underlying heuristic. In the case of feature selection in ML, the optimal subset of features is searched in the feature space based on some heuristic or performance measure. Generally, a metaheuristic algorithm consists of two phases: exploitation and exploration. In the exploitation or intensification phase, the algorithm explores the neighborhood of an already promising solution in the search space. However, during exploration (a.k.a., diversification), the algorithm tries to traverse the unvisited regions of the search space. Although they do not make any hypothesis on the mathematical properties of the objective function, they gradually develop it through a continuous learning process. Among the major advantages, the utilization of parameters and comparatively faster convergence to the solutions are crucial. In addition, metaheuristics are efficient and effective in obtaining global optimal values, resulting in global optimal features. Moreover, even with large datasets, they perform significantly well [dokeroglu2022comprehensive]. However, metaheuristics are approximate and usually non-deterministic and do not guarantee the optimal (or, best) solution [desale2015heuristic] like the exact algorithms (e.g., dynamic programming, branch and bound, branch and cut, linear programming, etc.). Still, they can provide near-optimal solutions in acceptable computing time (but the exact algorithms can not), which is highly essential for complex problems like detecting intrusions in dynamic environments. Interestingly, these techniques are mostly inspired by natural phenomena, including the instincts of living creatures. In addition to filter and wrapper techniques, a hybrid approach is another one that focuses on combining different aspects of existing feature selection methods [intro_3_7745366].
@@
Regarding attack or intrusion detection, traditional IDSs like statistical-driven (e.g., payload-based), rule-based, heuristics-based, etc. cannot detect the complex patterns of dynamic IoT systems. On the other hand, classical and deep machine learning models have been proven to generate notable results, even in heterogeneous environments such as IoT. The main purpose of using ML-based techniques is to handle large data sets and produce high accuracy, fast processing, and significant performance; thus enhancing security. However, they require high computational resources and a significant amount of time to achieve minor precision improvements [bharati2022machine]. Despite these improvements offered by machine learning, the era of big data and the increasing use of IoT introduce new problems with traditional centralized cloud-based data storage and processing systems. In particular, low throughput, high latency, and data privacy are the most serious issues [hua2023edge]. In addition, IoT devices contain sensitive and private data, such as financial or patient information. To address these issues, edge computing technology has become widely accepted, especially in the IoT context. In this technology, data are processed, stored, and computed closer to the location of devices. Consequently, not only the data transmission time, response time, and latency are reduced but also higher scalability and decentralization are achieved. Regarding IDSs, when machine learning models are trained in edge servers with large datasets, the computing power and the adequacy of energy support become crucial challenges since edge servers can hardly meet these requirements [yu2017survey].
@@
Although researchers always rely on utilizing optimization techniques to mitigate such problems and improve the effectiveness of ML-oriented IoT-IDSs, metaheuristics-based optimization has been a notable focus in recent years. Considering the outstanding facilities offered by these optimizers, they can play a pivotal role in designing IoT-IDSs. Specifically, these algorithms can be utilized not only to select optimal feature sets — before being trained by an ML-based classifier — but also to optimize the parameters (e.g., weights and biases) and hyperparameters (e.g., learning rate, number of neurons, layers volume, and amount of epochs) of the models during training. For these reasons, numerous recent works [1_kareem2022effective, 3_alkanhel2023network, 7_sanju2023enhancing, 97_10.1007/s11276-023-03435-0, 50_jovanovic2022feature] have employed them to select an optimal set of features; while others have used these techniques to tune parameters [35_baniasadi2022novel, 75_jothi2023wils] and hyperparameters [29_bahaa2022novel, 71_basheri2023quantum, 85_vijayan2024original] in ML-driven classifiers.
@@
The remainder of this paper proceeds as follows. Section 2 introduces the background, including the classification of intrusion detection systems, metaheuristics, and machine learning techniques. Section 3 provides the related surveys with their limitations, research gaps analysis, and differentiating aspects of our work. Then, in Section 4, we discuss our research objectives along with search strategy and data assessment. Next, Section 5 introduces the results of the systematic literature review, including the technical and extensive exploration of the existing relevant detection systems and many insightful findings. After that, in Section 6, our overall investigation with possible future challenges are summarized. Finally, Section 7 concludes our work. Along with them, an Appendix section at the end of the paper presents a performance tabulation of existing metaheuristics-based and ML-driven IoT-IDSs in tabulation form.
2 Background
@@
2.1 Intrusion Detection Methods and Techniques in IoT
@@
Regarding the IoT, most of the intrusion detection systems are network-based since attackers can seamlessly misuse the heterogeneous and dynamic characteristics of the IoT environment. In the literature, NIDS is classified into the following four methods: Signature-based Intrusion Detection System (SIDS), Anomaly-based Intrusion Detection System (AIDS), Specification-based Intrusion Detection System (SpIDS), and Hybrid Intrusion Detection System (HyIDS). In SIDS, an intrusion signature is checked with the previously known intrusion patterns, stored in the database, to find significant matching. It is also known as knowledge-based detection or misuse detection. AIDS is a dynamic intrusion detection approach that monitors the activity log of a system and reports anomalies whenever it observes any deviations from normal behaviors. This method provides the capabilities to detect not only known and unknown attacks but also any insider attacks. In SpIDS, a set of rules and thresholds are defined for network modules like nodes, protocols, firewalls, etc. Utilizing these specifications, the system detects intrusions while observing any discrepancies from the acceptable behaviors [mitchell2014survey]. On the other hand, Hybrid IDS incorporates the advantages of SIDS, AIDS, and SpIDS to detect both familiar and novel intrusions utilizing limited computational resources. The classification of NIDS is showed in Figure 1.
@@
Techniques. Considering the resource and energy-constrained characteristics, AIDS and HyIDS are the most appropriate and feasible methods in IoT [fu2017automata, midi2017kalis]. Various machine Learning-based, statistical-driven (like payload-based), rule-based, and heuristics-based techniques are used to escalate the training process for AIDS. Recently, metaheuristics and hybrid approaches integrating different algorithms have been developed in this field, especially for IoT. In the next two Sections, some well-established techniques used in IoT-IDS are discussed. Figure 2 illustrates the classification of the existing IoT-IDS techniques as a whole.
2.2 Metaheuristic Algorithms
@@
A metaheuristic is a general exploration (or diversification) method that can be applied to different problems in a similar way by visiting regions of the search space that are not already seen and evaluating candidate solutions. In general terms, metaheuristics are approximation algorithms that provide good or acceptable solutions within an acceptable computing time, that cannot be obtained with more specialized techniques, such as brute-force, linear programming, dynamic programming, randomization, quantum computation, exact algorithms, etc., but do not give formal guarantees about the quality of the solutions [chopard2018introduction]. The main difference between heuristic and metaheuristic is that a heuristic algorithm utilizes some specially designed functions to explore the solution space intelligently; whereas, a metaheuristic is an iterative generation process that directs a supporting heuristic to explore and exploit the search space efficiently [desale2015heuristic]. Moreover, heuristics can be applied to a specific problem; but, metaheuristics are more generalized and can be employed in the same way to many different problems. The metaheuristics algorithms can be classified into three groups as discussed below.
@@
Population-based metaheuristics: This type of metaheuristic utilizes global exploration and local exploitation ability for searching in global search space to discover new promising solutions and to refine the already discovered solutions. In this study, these algorithms are divided into three classes: (i) Swarm-based, (ii) Nature-based, and (iii) Evolutionary Algorithms (EAs). Swarm-based metaheuristics (a.k.a. swarm intelligence) are inspired by the social collective behavior of the birds, ants, bees, etc., where each animal (artificial agents) interacts with each other to achieve a particular goal in the environment [sherinov2018multi]. Ant colony optimization (ACO), particle swarm optimization (PSO), Artificial bee colony (ABC), etc. are some of the most popular swarm-based metaheuristics. The second category is nature-inspired optimization, such as gorilla troops optimizer (GTO), crow search algorithm (CrSA), reptile search algorithm (RSA), butterfly optimization algorithm (BOA), moth–flame optimization (MFO), biogeography-based optimization (BBO), intelligent water drop (IWD), etc. Particularly, these algorithms mimic the successful characteristics of complex natural processes, including distinct animal behaviors, biological systems, natural calamities, etc. Though swarm intelligence also relies on nature, it specifically focuses on decentralized systems and collaborative behaviors; whereas, nature-inspired metaheuristics encompass diverse elements of nature, including the behavior of an individual animal. Another well-established population-based optimizations are evolutionary algorithms, which are based on the process of natural evolution like survival, reproduction, and mutation. Specifically, there are three important components of an EA: parent selection, variation operators (recombination/crossover and mutation), and replacement (evolution). Genetic algorithms (GAs), evolutionary programming (EP), and differential evolution (DE) are the most popular EAs in the literature.
@@
Iterative-based metaheuristics: The second major group of metaheuristics is iterative-based. These algorithms are inspired by the laws of physics, mathematics, chemistry, or social human behavior. Particularly, physics-based metaheuristics are based on the concepts of physical laws and principles, for example, classical mechanics, thermodynamics, optics, etc. Gravitational search algorithm (GSA), simulated annealing (SA), multi-verse optimizer (MVO), etc. are the most popular physics-based optimization techniques. Similarly, math-based metaheuristics adopt mathematical concepts like number theory, geometry, and algebra, along with modern mathematics. Arithmetic optimization algorithm (AOA) [abualigah2021arithmetic] and sine cosine algorithm (SCA) [mirjalili2016sca] are prominent techniques in literature that are based on the arithmetic operators and sine/cosine mathematical functions, respectively. Other interesting iterative-based metaheuristics are inspired by human interaction, intelligence, learning processes, and experiences. Some of these algorithms are teaching-learning-based optimization (TLBO), human-guided search (HGS) [klau2010human], and harmony search (HS). The rest algorithms in this category are search-based, for instance, local search (LS), tabu search (TS), neighborhood search (NS), etc.
@@
2.3 Machine learning techniques

### v4

@@ Abstract
I Introduction
@@
The popularity of the Internet of Things devices has spread surprisingly in the last few years. Nowadays, it offers a scalable platform not only for industry, healthcare, and home applications but also for agriculture, vehicular settings, and ultra-sophisticated systems like drone technologies [4]. Alarmingly, this widespread adoption leads to unavoidable security issues as a negative side effect of close connectivity. Transferring susceptible information, such as personal data, patient data, and private business analysis makes such effects more severe and unimaginable. For example, millions of IoT and embedded devices (e.g., DVDs, printers, and IP cameras) were infected by a botnet in 2016, widely known as Mirai botnet[19]. Notably, the attack was spread to several nations and manufacturing organizations and affected around 65 thousand IoT devices within the first 20 hours. Another concerning incident occurred in 2020, where an ADT employee pled guilty to accessing the security cameras of 220 women over 9,600 times during four years [50]. The defendant routinely added his email address to customers’ “ADT Pulse” accounts and got real-time access to the video feeds from their homes for sexual gratification. In such cases, an IoT-IDS can be used to detect and report on unique visitors, unauthorized access, and malicious activities.
@@
Intrusion detection is one of the most crucial aspects of IoT security. An IDS typically identifies diverse attacks based on predefined rules or specific deviations from normal behavioral patterns. It can identify external and internal attacks on networks or computer systems, surpassing the performance of the traditional firewall. Typically, a firewall works on a set of specific rules, often based on IP addresses, port numbers, and protocols to check which packets are allowed to enter the network. Since firewalls filter packets relying on simple policies, they cannot detect internal or external attacks which require analyzing complex behavioral patterns. On the contrary, an IDS can identify malicious activities by observing the deviations from normal behaviors, which include both simple and complicated patterns. The entire IDS can be divided into two modules: feature engineering (mostly feature selection) and classification or detection. Feature selection aims to select a set of optimal features, discarding the least significant ones to efficiently and faster the classification process and reduce the computational overhead of the system. Feature selection methods can be categorized into three methods: filter, wrapper, and hybrid methods. In the filter method, all features are statistically examined and rated with the help of data (input and target variables). Then based on the rating, less ranked features or features rated below a specific threshold are eliminated before classification. Information gain and linear correlation coefficient [52] are well-known metrics used in filter methods. The wrapper techniques outperform filter methods by training and testing a machine-learning model using each subset of features, iteratively. Specifically, these methods rank subsets of features based on their prediction accuracy generated from the machine learning algorithms. However, they are more expensive and time-consuming than the filter approaches. Sequential forward selection, sequential backward selection, stepwise selection [79], hill climbing, etc. are popular wrapper methods. Though filter methods are proven simple, fast, and scalable, they consider feature dependencies to a certain extent, resulting in an inappropriate feature set. Besides, wrapper methods offer better feature sets than filters. But, they become much slower and computationally expensive for a large number of features [171].
@@
A group of wrapper methods fall into the category of optimization-based techniques, known as metaheuristics algorithms that overcome the drawbacks of the prior methods. A metaheuristic is a general exploration method that applies to optimize an underlying heuristic. In the case of feature selection in ML, the optimal subset of features is searched in the feature space based on some heuristic or performance measure. Generally, a metaheuristic algorithm consists of two phases: exploitation and exploration. In the exploitation or intensification phase, the algorithm explores the neighborhood of an already promising solution in the search space. However, during exploration (a.k.a., diversification), the algorithm tries to traverse the unvisited regions of the search space. Although they do not make any hypothesis on the mathematical properties of the objective function, they gradually develop it through a continuous learning process. Among the major advantages, the utilization of parameters and comparatively faster convergence to the solutions are crucial. In addition, metaheuristics are efficient and effective in obtaining global optimal values, resulting in global optimal features. Moreover, even with large datasets, they perform significantly well [49]. However, metaheuristics are approximate and usually non-deterministic and do not guarantee the optimal (or, best) solution [44] like the exact algorithms (e.g., dynamic programming, branch and bound, branch and cut, linear programming, etc.). Still, they can provide near-optimal solutions in acceptable computing time (but the exact algorithms can not), which is highly essential for complex problems like detecting intrusions in dynamic environments. Interestingly, these techniques are mostly inspired by natural phenomena, including the instincts of living creatures. In addition to filter and wrapper techniques, a hybrid approach is another one that focuses on combining different aspects of existing feature selection methods [53].
@@
Regarding attack or intrusion detection, traditional IDSs like statistical-driven (e.g., payload-based), rule-based, heuristics-based, etc. cannot detect the complex patterns of dynamic IoT systems. On the other hand, classical and deep machine learning models have been proven to generate notable results, even in heterogeneous environments such as IoT. The main purpose of using ML-based techniques is to handle large data sets and produce high accuracy, fast processing, and significant performance; thus enhancing security. However, they require high computational resources and a significant amount of time to achieve minor precision improvements [33]. Despite these improvements offered by machine learning, the era of big data and the increasing use of IoT introduce new problems with traditional centralized cloud-based data storage and processing systems. In particular, low throughput, high latency, and data privacy are the most serious issues [83]. In addition, IoT devices contain sensitive and private data, such as financial or patient information. To address these issues, edge computing technology has become widely accepted, especially in the IoT context. In this technology, data are processed, stored, and computed closer to the location of devices. Consequently, not only the data transmission time, response time, and latency are reduced but also higher scalability and decentralization are achieved. Regarding IDSs, when machine learning models are trained in edge servers with large datasets, the computing power and the adequacy of energy support become crucial challenges since edge servers can hardly meet these requirements [175].
@@
Although researchers always rely on utilizing optimization techniques to mitigate such problems and improve the effectiveness of ML-oriented IoT-IDSs, metaheuristics-based optimization has been a notable focus in recent years. Considering the outstanding facilities offered by these optimizers, they can play a pivotal role in designing IoT-IDSs. Specifically, these algorithms can be utilized not only to select optimal feature sets — before being trained by an ML-based classifier — but also to optimize the parameters (e.g., weights and biases) and hyperparameters (e.g., learning rate, number of neurons, layers volume, and amount of epochs) of the models during training. For these reasons, numerous recent works [91, 13, 150, 93, 89] have employed them to select an optimal set of features; while others have used these techniques to tune parameters [28, 88] and hyperparameters [26, 29, 170] in ML-driven classifiers.
@@
The remainder of this paper proceeds as follows. Section II introduces the background, including the classification of intrusion detection systems, metaheuristics, and machine learning techniques. Section III provides the related surveys with their limitations, research gaps analysis, and differentiating aspects of our work. Then, in Section IV, we discuss our research objectives along with search strategy and data assessment. Next, Section V introduces the results of the systematic literature review, including the technical and extensive exploration of the existing relevant detection systems and many insightful findings. After that, in Section VI, our overall investigation with possible future challenges are summarized. Finally, Section VII concludes our work. Along with them, an Appendix section at the end of the paper presents a performance tabulation of existing metaheuristics-based and ML-driven IoT-IDSs in tabulation form.
II Background
@@
II-A Intrusion Detection Methods and Techniques in IoT
@@
Regarding the IoT, most of the intrusion detection systems are network-based since attackers can seamlessly misuse the heterogeneous and dynamic characteristics of the IoT environment. In the literature, NIDS is classified into the following four methods: Signature-based Intrusion Detection System (SIDS), Anomaly-based Intrusion Detection System (AIDS), Specification-based Intrusion Detection System (SpIDS), and Hybrid Intrusion Detection System (HyIDS). In SIDS, an intrusion signature is checked with the previously known intrusion patterns, stored in the database, to find significant matching. It is also known as knowledge-based detection or misuse detection. AIDS is a dynamic intrusion detection approach that monitors the activity log of a system and reports anomalies whenever it observes any deviations from normal behaviors. This method provides the capabilities to detect not only known and unknown attacks but also any insider attacks. In SpIDS, a set of rules and thresholds are defined for network modules like nodes, protocols, firewalls, etc. Utilizing these specifications, the system detects intrusions while observing any discrepancies from the acceptable behaviors [121]. On the other hand, Hybrid IDS incorporates the advantages of SIDS, AIDS, and SpIDS to detect both familiar and novel intrusions utilizing limited computational resources. The classification of NIDS is showed in Figure 1.
@@
Techniques. Considering the resource and energy-constrained characteristics, AIDS and HyIDS are the most appropriate and feasible methods in IoT [64, 119]. Various machine Learning-based, statistical-driven (like payload-based), rule-based, and heuristics-based techniques are used to escalate the training process for AIDS. Recently, metaheuristics and hybrid approaches integrating different algorithms have been developed in this field, especially for IoT. In the next two Sections, some well-established techniques used in IoT-IDS are discussed. Figure 2 illustrates the classification of the existing IoT-IDS techniques as a whole.
II-B Metaheuristic Algorithms
@@
A metaheuristic is a general exploration (or diversification) method that can be applied to different problems in a similar way by visiting regions of the search space that are not already seen and evaluating candidate solutions. In general terms, metaheuristics are approximation algorithms that provide good or acceptable solutions within an acceptable computing time, that cannot be obtained with more specialized techniques, such as brute-force, linear programming, dynamic programming, randomization, quantum computation, exact algorithms, etc., but do not give formal guarantees about the quality of the solutions [38]. The main difference between heuristic and metaheuristic is that a heuristic algorithm utilizes some specially designed functions to explore the solution space intelligently; whereas, a metaheuristic is an iterative generation process that directs a supporting heuristic to explore and exploit the search space efficiently [44]. Moreover, heuristics can be applied to a specific problem; but, metaheuristics are more generalized and can be employed in the same way to many different problems. The metaheuristics algorithms can be classified into three groups as discussed below.
@@
Population-based metaheuristics: This type of metaheuristic utilizes global exploration and local exploitation ability for searching in global search space to discover new promising solutions and to refine the already discovered solutions. In this study, these algorithms are divided into three classes: (i) Swarm-based, (ii) Nature-based, and (iii) Evolutionary Algorithms (EAs). Swarm-based metaheuristics (a.k.a. swarm intelligence) are inspired by the social collective behavior of the birds, ants, bees, etc., where each animal (artificial agents) interacts with each other to achieve a particular goal in the environment [158]. Ant colony optimization (ACO), particle swarm optimization (PSO), Artificial bee colony (ABC), etc. are some of the most popular swarm-based metaheuristics. The second category is nature-inspired optimization, such as gorilla troops optimizer (GTO), crow search algorithm (CrSA), reptile search algorithm (RSA), butterfly optimization algorithm (BOA), moth–flame optimization (MFO), biogeography-based optimization (BBO), intelligent water drop (IWD), etc. Particularly, these algorithms mimic the successful characteristics of complex natural processes, including distinct animal behaviors, biological systems, natural calamities, etc. Though swarm intelligence also relies on nature, it specifically focuses on decentralized systems and collaborative behaviors; whereas, nature-inspired metaheuristics encompass diverse elements of nature, including the behavior of an individual animal. Another well-established population-based optimizations are evolutionary algorithms, which are based on the process of natural evolution like survival, reproduction, and mutation. Specifically, there are three important components of an EA: parent selection, variation operators (recombination/crossover and mutation), and replacement (evolution). Genetic algorithms (GAs), evolutionary programming (EP), and differential evolution (DE) are the most popular EAs in the literature.
@@
Iterative-based metaheuristics: The second major group of metaheuristics is iterative-based. These algorithms are inspired by the laws of physics, mathematics, chemistry, or social human behavior. Particularly, physics-based metaheuristics are based on the concepts of physical laws and principles, for example, classical mechanics, thermodynamics, optics, etc. Gravitational search algorithm (GSA), simulated annealing (SA), multi-verse optimizer (MVO), etc. are the most popular physics-based optimization techniques. Similarly, math-based metaheuristics adopt mathematical concepts like number theory, geometry, and algebra, along with modern mathematics. Arithmetic optimization algorithm (AOA) [3] and sine cosine algorithm (SCA) [120] are prominent techniques in literature that are based on the arithmetic operators and sine/cosine mathematical functions, respectively. Other interesting iterative-based metaheuristics are inspired by human interaction, intelligence, learning processes, and experiences. Some of these algorithms are teaching-learning-based optimization (TLBO), human-guided search (HGS) [100], and harmony search (HS). The rest algorithms in this category are search-based, for instance, local search (LS), tabu search (TS), neighborhood search (NS), etc.
@@
II-C Machine learning techniques